demolisher1224:

plz help me im struggleing what is the solution to the system below? 2x+y=3 4x+2y=6 (2,2) (2,-2) no solution infinite solutions on the line

dude:

Uh you can solve this using substitution, do you know how to do that?

demolisher1224:

dude i dont pay attention in my live lessons because i work a job 24 7

dude:

D: sorry Well I'l walk you through it What is the opposite of adding 2x?

demolisher1224:

subract 2x from both sides right

dude:

Right (We're doing this on the first equation 2x+y=3)

demolisher1224:

2x+y=3 -2x+y=-2x am i right so far

dude:

Close  Don't forget the 3 and the 2x on the left side is crossed out

demolisher1224:

so now we have to either get rid of the x or we keep as is

dude:

Keep it as is, we know what y equals, that's all we wanted for now Then we have to substitute the equation we just got into the second equation y=-2x+3 into -> 4x+2y=6 Do you know how to do that?

demolisher1224:

no i dont remember

mikimoff18:

The answer is the same for both x=(-1/2)y+3/2 y=-2x+3

dude:

 We know what y is, we just insert it into the second equation

demolisher1224:

so im supost to subtract it now and i will get my answer

dude:

Well, almost, you first want to distribute 2(-2x+3)

demolisher1224:

-4x+6

dude:

Good good 4x-4x+6=6 Now we combine the x's

dude:

We should get 6=6 Because they are the same this means that there are infinitely many solutions

demolisher1224:

ok thx